OpenStudy (anonymous):

if A=90DEGREE SHOW THAT 2=b+c where r is the inradius

OpenStudy (anonymous):

Is it an inequality or a bracket

OpenStudy (anonymous):

bracket

OpenStudy (anonymous):

As sin 90=1 so 2R=2RsinA=a and r=s-a*tana/2 so tan A/2=1 so 2(r+r)=a+2[s-a]=b+c-a+a=b+c

OpenStudy (anonymous):

2(R+r)

OpenStudy (anonymous):

r=[s-a]tan A/2